OpenStudy (anonymous):

There are 3 large marbles, 2 medium marbles and 5 small marbles in a bag. If one of the marbles is chosen randomly, what is the probability that a small marble is chosen?

OpenStudy (ksaimouli):

abhoooooooooooooooo chalanaudhi ra

OpenStudy (ksaimouli):

2

OpenStudy (ksaimouli):

add 3+2+5 they asked proba so divde 5/10=2

OpenStudy (anonymous):

The total number of marbles are 10. So the odds of any one marble being chosen are 1/10. 5 of the marbles are small. So there's 5/10 = 1/2 chance of a small marble being chosen.

OpenStudy (anonymous):

MORON

OpenStudy (anonymous):

... thanks.